OpenStudy (anonymous):

Roberta deposited $3,116.92 into a savings account with an interest rate of 3.9% compounded quarterly. About how long will it take for the account to be worth $7,000?

OpenStudy (anonymous):

Options: a) 14 years, 5 months b) 20 years, 10 months c) 21 years, 2 months d) 29 years, 11 months

OpenStudy (nurali):

Use the Formula, A = P(1 + i/n)^(nt) A is the eventual amount, P is the Principal amount, i is the interest expressed as a decimal, n is the number of times the interest is compounded per time period and t is the number of time periods. 7000 = 3116.92(1 + 0.039/4)^4t 7000= 3116.92(1.00975)^4t 7000/3116.92 = 1.00975^4t 2.24581= 1.00975^4t log on both side 4tLog 1.00975 = Log 2.24581 4t ≈ 83.385

OpenStudy (nurali):

can you solve for t now?

OpenStudy (anonymous):

I dont understand the log thingy... :/

OpenStudy (anonymous):

the answer will be 20 years and 10month.... but i want to understand how you got to: 4t ≈ 83.385 (like, how do you solve the log, and why u used it? )